OpenStudy (anonymous):

At the beginning of an experiment, a scientist has 112 grams of radioactive goo. After 225 minutes, her sample has decayed to 14 grams. What is the half-life of the goo in minutes? Find a formula for , the amount of goo remaining at time.? How many grams of goo will remain after 26 minutes?

OpenStudy (amistre64):

14 = 112e^(225r) 14 ---- = e^(225r) 112 ln(1/8) = 225r r = ln(1/8)/225 ................................... y = 112e^(ln(1/8)/225)t lol

OpenStudy (amistre64):

112/2 = 56 56 ---- = 1/2 ..... im getting dumber by the minute lol 112 ln(1/2) ------ = halftime rate

OpenStudy (amistre64):

half-life = 75 minutes

OpenStudy (amistre64):

and im thinking 8' 2/3 after 26 minutes

OpenStudy (amistre64):

but that last part is prolly wrong

OpenStudy (anonymous):

lol ur fine i thank you for trying and helping:)